In a 2 digit number , the units digit is twice the tens digit . If 27 is added to the number the digit inter-change their places . Find the number ? Pls help me i will be thankful
Answer by mszlmb(115)

the unit's digit is twice the ten's digit; so the unit's digit can only be 2,4,6, or 8. [it cannot be 1 because 1/2 is .5 and you can't have a half in the ten's digit, can't be 3,5,7,9 because they're odd and require a half also, can't be zero unless ten's is zero but it's a 2 digit number so that can't be]
Accordingly, the ten's digit can only be 1,2,3,4. Can go no higher because twice 5 is more than 9 (highest 1 dgt number) and so forth.
the choices are 12,24,36, and 48. No more, as we have deduced, would fit the scenerio. This makes the problem relatively easier, we can use guess and check:
12+27=39 nope
24+27=51 nope
36+27=63 YES
48+27=75 nope
..THERE'S YOUR ANSWER!